# Revisiting Global Income Convergence in the 21st Century

### v1

@@ 5 Accounting for Decline in Income Dispersion
Here, the second and third terms capture the role of inputs, while the residual term ensures that the accounting identity holds. Appendix A provides a detailed explanation of how each component is computed.
I report the results from the accounting exercise for three percentile ratios-P90/P10, P90/P50, and P50/P10-covering countries outside Sub-Saharan Africa for the years 2000-2019, and 1980-2019. These ratios capture the overall trend in convergence outside Sub-Saharan Africa across the income distribution.
Table 3: Accounting for Income Convergence outside SSA
Percentile Ratio $\Delta_{t}\log(\frac{y(p_{n})}{y(p_{0})})$ Cont. Cont. Cont.
(%) A (%) $\frac{K}{Y}$ (%) $h$ (%)
A: 1980-2000
P90/P10 8 487 -225 -175
P90/P50 18 194 -67 -22
P50/P10 -10 -40 60 90
B: 2000-2019
P90/P10 -54 44 43 11
P90/P50 -34 35 44 26
P50/P10 -20 65 40 -15
C: 1980-2019
P90/P10 -46 -33 89 43
P90/P50 -16 -144 169 81
P50/P10 -30 30 47 20
Table 3 presents the results of the accounting exercise, showing changes in income percentile ratios over three periods: 1980-2000, 2000-2019, and 1980-2019. Positive values in the changes of income percentile ratios signify income divergence, indicating that income inequality between countries has increased, while negative values denote income convergence, meaning that income inequality has decreased. The table breaks down the contributions to these changes from three factors: total factor productivity (TFP) differences, capital-output ratios, and human capital differences. Each row of the table details these contributions across three key percentile ratios-P90/P10, P90/P50, and P50/P10-which represent disparities between the rich (90th percentile) and the poor (10th percentile), the rich and the middle-income group (50th percentile), and the middle-income group and the poor, respectively. The sign of each percentage contribution indicates the direction of its contribution relative to the overall change: a positive percentage means the factor’s change aligns with the total change, while a negative percentage means it opposes the total change. For example, during a period of income divergence (a positive change in the percentile ratio), a positive contribution from TFP differences implies that TFP is also diverged during the same period. Similarly, if the percentile ratio is converging (a negative change), a positive contribution from TFP differences indicates that TFP is also converged during the same period.
Panel A of Table 3 shows that between 1980 and 2000, the P90/P10 and P90/P50 income ratios diverged, while the P50/P10 ratio converged. Although the capital-output ratio and human capital converged between rich and poor economies and between rich and middle-income economies, TFP diverged among these groups, leading to overall income divergence. Notably, the income divergence would have been even greater if inputs had not converged. In contrast, the income difference between middle-income and poor economies converged during this period. The convergence of inputs accounts entirely for the income convergence between middle-income and poor economies, as TFP diverged between these groups during this time.
Panel B reveals that, unlike the 1980-2000 period, income differences converged across the income distribution during 2000-2019. This shift was due to TFP diverging in the earlier period to converging in the later period. More importantly, the reduction in the gaps in the capital-output ratio significantly contributed to the narrowing of income differences during 2000-2019. Specifically, the P90/P10 income ratio declined by 54%, with the decline in the capital-output ratio accounting for 43% of this reduction. Similarly, the P90/P50 ratio declined by 34%, and the P50/P10 ratio declined by 20%. The decline in the capital-output ratio explains 44% of the reduction in P90/P50 differences and 40% of the reduction in P50/P10 differences. These findings highlight the critical role of physical capital in driving convergence during this period. When changes in the capital-output ratio are combined with improvements in human capital, the total contribution of inputs becomes even more pronounced: inputs account for 52% of the reduction in P90/P10 disparities, 65% of the reduction in P90/P50 disparities, and 35% of the reduction in P50/P10 disparities. The overall decline in input dispersion accounts for more than half of the reduction in output disparities between the richest (P90) and poorest (P10) economies, as well as between the richest and middle-income (P50) economies.
Panel C presents the findings for the extended period from 1980 to 2019. During this timeframe, convergence in the capital-output ratio was the primary driver of narrowing income disparities. The P90/P10, P90/P50, and P50/P10 income ratios declined by 46%, 16%, and 30%, respectively. The reduction in the capital-output ratio accounts for 89% of the decrease in P90/P10 income differences, while TFP differences contributed negatively with -33%, indicating divergence. For the P90/P50 ratio, the capital-output ratio explains a striking 169% of the reduction, offsetting the negative contribution from the TFP divergence of -144%. For the P50/P10 ratio, the capital-output ratio accounts for 47% of the convergence, with TFP contributing 30%. When changes in the capital-output ratio are combined with improvements in human capital, inputs account for 132% of the reduction in P90/P10 disparities, 250% of the reduction in P90/P50 disparities, and 77% of the reduction in P50/P10 disparities. These results imply that, had it not been for the significant convergence in the inputs, especially the capital-output ratio, the overall income convergence between the richest and poorer countries would have been substantially lower or potentially nonexistent due to TFP divergence during this period.
5.1 Accounting with constant capital income share
Table 4: Accounting for Income Convergence outside SSA ($\alpha=1/3$)
Percentile Ratio $\Delta_{t}\log(\frac{y(p_{n})}{y(p_{0})})$ Cont. Cont. Cont.
(%) A (%) $\frac{K}{Y}$ (%) $h$ (%)
A: 1980-2000
P90/P10 8 412 -137 -175
P90/P50 18 178 -44 -22
P50/P10 -10 -10 30 90
B: 2000-2019
P90/P10 -54 65 24 11
P90/P50 -34 53 24 26
P50/P10 -20 85 25 -15
C: 1980-2019
P90/P10 -46 4 52 43
P90/P50 -16 -88 100 81
P50/P10 -30 53 27 20
Table 4 presents the main accounting result assuming a constant capital share of one-third. Panel A of Table 4 shows that between 1980 and 2000, the divergence of TFP among rich and poor economies, as well as between poor and middle income economies still accounts for all the divergence in income ratios. This suggests that for studies focused on this time period, such as [14], the assumption of a constant capital share does not make much of a difference. Panel B of Table 4, shows that between 2000 and 2019, TFP accounts for the majority of convergence among rich and poor, middle and poor, and rich and middle income economies. This contrasts with the results in Panel B of Table 3, where inputs accounts for majority of the convergence. Therefore, for the period post-2000, the choice of capital income share becomes crucial. When heterogeneity of capital income shares is incorporated into the growth accounting exercise, capital alone accounts for 43%, 44%, and 40% of the convergence in P90/P10, P90/P50, and P50/P10 ratios between 2000 and 2019. In comparison, with the assumption of a constant capital income share of one-third, the convergence of capital accounts for only 24%, 24%, and 25% of the convergence, respectively. Hence, not incorporating heterogeneity in capital income shares during this time period would lead to the incorrect conclusion that TFP remains the dominant factor driving growth and convergence. Panel C of Table 4 shows that for the extended period from 1980 to 2019, with a constant one-third capital income share, convergence of capital accounts for 52%, and 27% of the convergence in P90/P10 and P50/P10 ratios, respectively, compared to 89% and 47% when heterogeneity of capital income shares is included. Capital still accounts for all the convergence in the P90/P50 ratio in both cases. Therefore, even in the long run, the assumption of a constant capital income share significantly reduces the role of capital in driving income convergence between rich and poor, and between middle-income and poor economies.
Overall, the findings indicate that the convergence of inputs-particularly physical capital-accounts for all the income convergence between rich and poor economies during 1980-2019, and more than half of the convergence during 2000-2019 outside Sub-Saharan Africa. Moreover, the choice of capital income share is important. Assuming a constant capital share of one-third reduces the role of capital in driving convergence by approximately half.
@@ 6 Measurement of Inputs and Capital Share of Income
In this paper, the capital-output ratio is calculated as the ratio of real capital stock ($rnna$) to real GDP ($rgdpna$) at constant national prices from the Penn World Tables (PWT 10.01).
The capital stock series in the PWT is constructed using the Perpetual Inventory Method (PIM). The stock of capital of type $a$ at time $t$ is computed as the accumulated type-specific investment ($I_{at}$) net of depreciation ($\delta_{a}$):
@@
The aggregate capital stock is simply the sum of asset specific capital stocks, $K_{t}=\sum_{a}K_{at}$.
The PIM requires an assumption about the initial capital stock ($K_{0}$), i.e., the capital stock in the first year when investment data becomes available. A common approach in the literature is to use the steady-state relation from the neoclassical growth model:
@@
Measurement of Capital Share of Income
@@
where labsh represents the share of labor income in GDP. One significant challenge in measuring the labor share is that the labor income of self-employed workers is not directly observable. [9] discusses various methods for estimating the labor compensation of self-employed workers. Two common adjustments rely on mixed-income data from national accounts. Mixed income refers to the total income earned by self-employed workers, encompassing both labor and capital income. The first adjustment allocates all mixed income to labor, while the second adjustment allocates mixed income between labor and capital in the same proportion as the rest of the economy. In cases where mixed income data is unavailable, a third adjustment uses data on the number of employees and self-employed workers, assuming that the self-employed earn the same average wage as employees. PWT10.01 introduces a fourth adjustment that adds the value-added from agriculture to labor compensation of employees. Whenever mixed income data is available, PWT10.01 applies the second adjustment. In the absence of mixed-income data, I use the minimum of the third and fourth adjustments.
Figure 1: Variation of alpha over income percentile.
Figure 1 illustrates the capital share of income across countries at different percentiles of the income distribution for various years. For the sub-sample of countries included in this analysis, PWT shows an average capital share of income of 0.46, which significantly exceeds the commonly assumed value of 1/3 in the literature. Moreover, the capital share of income follows an inverse U-shaped pattern across the income distribution, with middle-income countries having the highest capital income share, followed by poor countries, and high-income countries having the lowest capital income share. Only the richest countries have capital income shares close to 1/3, suggesting that using a constant capital income share may not be appropriate for countries at other points in the income distribution. The decomposition approach in the current paper accommodates this heterogeneity, allowing the capital income share to vary along the income distribution. Additionally,, over time, countries across the income distribution have experienced an increase in the capital share of income, which indicates a corresponding decline in labor share. The recent work by [13] examines this global decline in labor share and attributes much of the trend to cheaper information and communication technology (ICT) capital. The growing importance of the capital share of income reinforces the hypothesis that the convergence of capital has played a substantial role in the convergence of output post-2000.
@@ 7 Conclusion
This study revisited the dynamics of cross-country income convergence using the most recent data from the Penn World Tables (PWT 10.01). Roughly 70% of cross-country empirical work relies on the Penn World Tables [11]. The updated PWT data no longer support some of the basic growth accounting facts that underpinned earlier studies. Contrary to the traditional view that emphasizes non-convergence and the limited role of factor inputs, the analysis reveals a significant shift in global income patterns since 2000. There has been unconditional convergence, particularly outside Sub-Saharan Africa. More importantly, accumulation of inputs, particularly physical capital, has significantly driven income convergence outside Sub-Saharan Africa.
@@ A.1 Generalized Decomposition
Let $p$ denote percentile and output be some arbitrary function if inputs, TFP parametrized by $p$.
$y(p)=F(A(p),k/y(p),h(p)).$
Then,
@@
$\log(y(p_{n}))-\log(y(p_{0}))$ $=\int_{p_{0}}^{p_{n}}[\frac{d\log F(.)}{d\log A(p)}\frac{d\log A(p)}{dp}+\frac{d\log F(.)}{d\log k/y(p)}\frac{d\log k/y(p)}{dp}+\frac{d\log F(.)}{d\log h(p)}\frac{d\log h(p)}{dp}]$
@@
where
$\mathcal{E}_{A}(p)$ $=\frac{d\log F}{d\log A}\frac{d\log A}{dp}$
$\mathcal{E}_{k/y}(p)$ $=\frac{d\log F}{d\log(k/y)}\frac{d\log(k/y)}{dp}$
$\mathcal{E}_{h}(p)$ $=\frac{d\log F}{d\log(h)}\frac{d\log(h)}{dp}$
Now suppose that the output production has the following functional form:
$y(p)$ $=A(p)(k(p))^{\alpha(p)}(h(p))^{1-\alpha(p)}$
$=A(p)^{\frac{1}{1-\alpha(p)}}(k/y(p))^{\frac{\alpha(p)}{1-\alpha(p)}}h(p)$
In this case:
$\mathcal{E}_{A}(p)$ $=\frac{1}{1-\alpha(p)}\frac{d\log A}{dp}$
$\mathcal{E}_{k/y}(p)$ $=\frac{\alpha(p)}{1-\alpha(p)}\frac{d\log(k/y)}{dp}$
$\mathcal{E}_{h}(p)$ $=\frac{d\log(h)}{dp}$
@@
Note that if the elasticity parameter is constant along income distribution ( $\alpha(p)=\alpha$), the the generalized decomposition reduces to the usual Hall-Jones decomposition.

### v2

@@ 4 Accounting for Decline in Income Dispersion
where $\alpha(p)$ denotes the capital share of income at percentile $p$. Here, the first term captures the contribution of TFP, the second term captures the contribution of capital-output ratio, and the third term captures the contribution of human capital. Appendix A provides the full derivations of each elasticity under CRS.
I report the results from the decomposition exercise for the percentile ratio-P90/P10- covering countries outside Sub-Saharan Africa for the time periods: 1980-2000, 2000-2019. The decomposition results for P90/P50, P50/P10, as well as for full sample of countries are reported in Appendix. [^7] I refer P90, P50, P10 as rich, middle, poor economies respectively.
Decomposition Results: Convergence outside SSA
Figure 1: Decomposition Exercise (excluding SSA): P90/P10
Figure 1 illustrates the decomposition of the P90/P10 income ratio over two distinct time periods: 1980-2000 (blue bars), and 2000-2019 (red bars). The vertical axis shows the change in values for each period, while the horizontal axis shows the overall change in the log P90/P10 ratio followed by the contributions of TFP, the capital-output ratio, and human capital to that change, as derived in equation (2).
Convergence in both physical and human capital inputs was the principal force shaping the evolution of the P90/P10 income ratio. Between 1980 and 2000, the P90/P10 income ratio increased by 8 percentage points, reflecting a period of divergence between rich and poor countries. Before 2000, TFP divergence alone would have driven the ratio up by roughly 0.39, but simultaneous convergence in physical capital and human capital reduced the ratio by about 0.17 and 0.14, respectively.
From 2000 to 2019 the pattern reversed. The log P90/P10 ratio fell by $0.54$. Convergence in the capital-output ratio contributed $0.24$ of this reduction (about 44%), while human-capital convergence contributed $0.06$ (10%). During this interval TFP shifted from divergence to convergence, contributing $0.25$ (46%) to overall convergence. The shift in total factor productivity (TFP) dynamics—from divergence before 2000 to convergence thereafter—played a pivotal role in driving post-2000 income convergence. In the absence of TFP convergence, the extent of overall convergence during this period would have been minimal. Nevertheless, in quantitative terms, the contribution of TFP to post-2000 convergence remains smaller than that of input accumulation, accounting for 46% compared to 54%. This contrast is even more pronounced when examining the earlier period from 1980 to 2000.
Note that the overall change in the P90/P10 ratio from 1980 to 2019 is the sum of the changes over 1980-2000 and 2000-2019, as expressed by the additive property of logarithms. Over the full horizon 1980-2019 the log P90/P10 ratio declined by $0.46$. Cumulatively, the capital-output ratio accounted for $0.41$ of the decline (89%) and human capital for $0.19$ (41%), whereas persistent TFP divergence added back roughly $0.14$. Despite the continued divergence in TFP, there was substantial convergence in income over the period 1980-2019. This implies that convergence in physical and human capital inputs fully accounts for the observed income convergence during this period.
The Appendix reports analogous decompositions for the P90/P50 and P50/P10 ratios, respectively. For the P90/P50 ratio input convergence accounts for 66% of the decline observed in 2000-2019 and explains the entire convergence over the full 1980-2019 horizon. In the case of the P50/P10 ratio inputs contribute less—32% of the convergence during 2000-2019 and 70% over 1980-2019. In both decompositions the behavior of total factor productivity mirrors that documented for the P90/P10 ratio, diverging in 1980-2000 and converging in 2000-2019. Taken together, the evidence indicates that poorer economies outside Sub-Saharan Africa are catching up not only with rich but also with middle-income countries. This long-run convergence—particularly after 2000, and more importantly over the full period from 1980 to 2019—is driven primarily by the convergence of physical and human capital inputs.
Decomposition Results: Stagnation of SSA
Figure 2: Decomposition Exercise (full sample): P90/P10
Figure 2 presents the decomposition results for the full sample of countries, including those in Sub-Saharan Africa (SSA). It is important to note that, as of 2000, economies below the P10 income level were predominantly from Sub-Saharan Africa. Therefore, this decomposition provides insights into the stagnation of poor economies in SSA relative to those outside the region. Similar to poor countries outside SSA, economies in SSA experienced divergence in income during 1980-2000 and convergence toward richer countries after 2000. However, the degree of divergence before 2000 was substantially larger within SSA (0.54 compared to 0.08), and the extent of convergence since 2000 was correspondingly much smaller (0.27 compared to 0.54). The larger divergence before 2000 for SSA economies is predominantly driven by more pronounced divergence in total factor productivity (TFP): 0.71 compared to 0.39 for non-SSA countries. Lower convergence in human capital also contributed: 0.03 for SSA compared to 0.14 outside SSA—but its role was relatively modest in comparison to TFP. Convergence in physical capital, by contrast, was broadly similar across the two groups, at 0.13 for SSA and 0.17 for non-SSA economies.
In the post-2000 period, convergence in TFP (0.22 compared to 0.25) and in human capital (0.05 compared to 0.06) were of similar magnitudes between SSA and non-SSA poor economies. The key distinction lies in physical capital: the absence of convergence in physical capital among SSA economies entirely accounts for the lower extent of output convergence relative to poor countries outside the region.
Unlike poor economies outside SSA, SSA did not exhibit any meaningful convergence in the capital-output ratio post-2000, and only negligible convergence in human capital. While TFP did converge post-2000, the lack of convergence in inputs meant that the forces of convergence were insufficient to offset the TFP divergence that occurred prior to 2000. As a result, between 1980 and 2019, SSA economies fell further behind the rich economies due to TFP divergence in the early period and the absence of input convergence across the full time span. Further evidence of the lack of input convergence, particularly in capital, is presented in Figure 3, which illustrates the mean capital-output ratio, weighted by population size, across different regions. While the capital-output ratio has continued to rise in all regions outside Sub-Saharan Africa, SSA is the only region to exhibit a decline in its capital-output ratio post-2000. This decline likely reflects the region’s struggles with capital accumulation—challenges that may be linked to political instability, economic uncertainty, and capital flight, as discussed in [16].
Figure 3: Capital-Output Ratio across regions
Decomposition with constant capital income share
Figure 4: Variation of alpha over income percentile.
Figure 4 illustrates the capital share of income across countries at different percentiles of the income distribution for various years. For the full sample of countries included in this analysis, PWT shows an average capital share of income of 0.46, which significantly exceeds the commonly assumed value of 1/3 in the literature. Moreover, the capital share of income follows an inverse U-shaped pattern across the income distribution, with middle-income countries having the highest capital income share, followed by poor countries, and high-income countries having the lowest capital income share. Only the richest countries have capital income shares close to 1/3, suggesting that using a constant capital income share may not be appropriate for countries at other points in the income distribution. The decomposition approach in the current paper accommodates this heterogeneity, allowing the capital income share to vary along the income distribution. Additionally, over time, countries across the income distribution have experienced an increase in the capital share of income, which indicates a corresponding decline in labor share. The recent work by [12] examines this global decline in labor share and attributes much of the trend to cheaper information and communication technology (ICT) capital. Because capital-income shares differ systematically across the income distribution, any growth-accounting exercise must incorporate this heterogeneity. Figure 5 reports-for comparison—the results obtained when a constant capital share of one-third is imposed.
Figure 5: Decomposition Exercise (excluding SSA, alpha=1/3): P90/P10
During the 1980-2000 period cross-country income divergence is still driven overwhelmingly by the widening gap in total factor productivity, with the convergence of physical and human capital acting in the opposite direction and therefore dampening the TFP effect. Because TFP remains the dominant force in the accounting, studies that concentrate on the pre-2000 era—such as [13]—are not materially affected by whether one allows the capital share of income to vary across countries or, instead, fixes it at the conventional one-third benchmark. However, for the post-2000 period, the choice of capital income share becomes important. With a constant one-third share, the capital-output ratio accounts for only 23% of the decline in the P90/P10 ratio, compared to 44% when heterogeneity in capital income shares is incorporated. This difference implies that ignoring heterogeneity may lead to the erroneous conclusion that TFP remains the dominant factor in driving growth and convergence. Over the full period from 1980 to 2019, the constant one-third assumption results in capital convergence explaining 52% of the overall convergence, as opposed to 89% when capital income shares vary by income. Thus, even in the long run, assuming a constant capital share significantly underestimates the role of capital in driving income convergence between rich and poor economies.
The Appendix further presents the decomposition results for the P90/P50 and P50/P10 ratios under the constant one-third assumption, showing a quantitatively reduced role of capital in driving convergence compared to the case when capital income shares are allowed to vary. The results remains unchanged when looking at full sample of countries.
@@ 5 Measurement of Inputs and Capital Share of Income
In this paper, the capital-output ratio is calculated as the ratio of real capital stock ($rnna$) to real GDP ($rgdpna$) at constant national prices from the Penn World Tables (PWT 10.01). The capital stock series in the PWT is constructed using the Perpetual Inventory Method (PIM). The stock of capital of type $a$ at time $t$ is computed as the accumulated type-specific investment ($I_{at}$) net of depreciation ($\delta_{a}$):
@@
The aggregate capital stock is simply the sum of asset specific capital stocks, $K_{t}=\sum_{a}K_{at}$. The PIM requires an assumption about the initial capital stock ($K_{0}$), i.e., the capital stock in the first year when investment data becomes available. A common approach in the literature is to use the steady-state relation from the neoclassical growth model:
@@
Capital Share of Income
@@
where labsh represents the share of labor income in GDP. One significant challenge in measuring the labor share is that the labor income of self-employed workers is not directly observable. [8] discusses various methods for estimating the labor compensation of self-employed workers. Two common adjustments rely on mixed-income data from national accounts. Mixed income refers to the total income earned by self-employed workers, encompassing both labor and capital income. The first adjustment allocates all mixed income to labor, while the second adjustment allocates mixed income between labor and capital in the same proportion as the rest of the economy. In cases where mixed income data is unavailable, a third adjustment uses data on the number of employees and self-employed workers, assuming that the self-employed earn the same average wage as employees. PWT10.01 introduces a fourth adjustment that adds the value-added from agriculture to labor compensation of employees. Whenever mixed income data is available, PWT10.01 applies the second adjustment. In the absence of mixed-income data, PWT10.01 uses the minimum of the third and fourth adjustments.
@@ 6 Conclusion
This study revisited the dynamics of cross-country income convergence using the most recent data from the Penn World Tables (PWT 10.01).999 Roughly 70% of cross-country empirical work relies on the Penn World Tables [10]. Contrary to the traditional view that emphasizes non-convergence and the limited role of factor inputs, the analysis reveals a significant shift in global income patterns since 2000. There has been unconditional convergence, particularly outside Sub-Saharan Africa. More importantly, accumulation of inputs, particularly physical capital, has significantly driven income convergence outside Sub-Saharan Africa.
@@
[…]
Acknowledgements
I am grateful to Anmol Bhandari, Alessandra Fogli, Hannes Malmberg, Todd Schoellman, and Kjetil Storesletten for their advice and support. I would also like to thank Honey Batra, Francisco Bullano, G V A Dharanan, Jason Hall, Sang Min Lee, Ruyan Liao, Martin Garcia Vazquez, and Teerat Wongrattanapiboon for useful comments.
@@ A.1 Generalized Decomposition
Let $p$ denote the percentile index, and consider the per-worker output function
$y(p)=F\bigl{(}A(p),\,k(p),\,h(p)\bigr{)}=F\bigl{(}A(p),\,(k(p)/y(p))\,y(p),\,h(p)\bigr{)},$ (3)
where $\kappa(p)=k(p)/y(p)$. Total differentiation of (3) with respect to $p$ gives
$\frac{dy}{dp}$ $=F_{A}\bigl{(}A,\kappa y,h\bigr{)}\,\frac{dA}{dp}+F_{k}\bigl{(}A,\kappa y,h\bigr{)}\,\frac{d(\kappa y)}{dp}+F_{h}\bigl{(}A,\kappa y,h\bigr{)}\,\frac{dh}{dp},$ (4)
$\frac{d(\kappa y)}{dp}$ $=\kappa\,\frac{dy}{dp}+y\,\frac{d\kappa}{dp}.$ (5)
Rearranging yields
$\bigl{[}1-F_{k}(A,\kappa y,h)\,\kappa\bigr{]}\,\frac{dy}{dp}=F_{A}(A,\kappa y,h)\,\frac{dA}{dp}+F_{k}(A,\kappa y,h)\,y\,\frac{d\kappa}{dp}+F_{h}(A,\kappa y,h)\,\frac{dh}{dp}.$ (6)
Dividing by $y$ gives the differential form
$\frac{d\ln y}{dp}=\underbrace{\frac{F_{A}}{y\,\bigl{(}1-F_{k}\,\kappa\bigr{)}}\,\frac{dA}{dp}}_{\mathcal{E}_{A}(p)}+\underbrace{\frac{F_{k}\,\kappa}{1-F_{k}\,\kappa}\,\frac{d\ln\kappa}{dp}}_{\mathcal{E}_{k/y}(p)}+\underbrace{\frac{F_{h}\,h}{y\,\bigl{(}1-F_{k}\,\kappa\bigr{)}}\,\frac{d\ln h}{dp}}_{\mathcal{E}_{h}(p)}.$ (7)
Integration from $p_{0}$ to $p_{n}$ yields the generalized decomposition
@@
For a CRS technology in $(k,h)$, Euler’s theorem implies
$F_{k}(A,k,h)\,k+F_{h}(A,k,h)\,h=y,\quad\alpha(p)=\frac{F_{k}(A,k,h)\,k}{y}.$ (8)
Hence
$F_{k}(A,k,h)\,\kappa=\alpha(p),\quad\frac{F_{h}(A,k,h)\,h}{y}=1-\alpha(p).$
Using these identities, the three elasticities in (7) become
$\mathcal{E}_{A}(p)$ $=\frac{F_{A}}{y\,(1-\alpha(p))}\,\frac{dA}{dp}=\frac{1}{1-\alpha(p)}\,\frac{d\ln A(p)}{dp},$
$\mathcal{E}_{k/y}(p)$ $=\frac{\alpha(p)}{1-\alpha(p)}\,\frac{d\ln(k/y)(p)}{dp},$
$\mathcal{E}_{h}(p)$ $=\frac{1-\alpha(p)}{1-\alpha(p)}\,\frac{d\ln h(p)}{dp}=\frac{d\ln h(p)}{dp}.$
Substituting back into the integrated form yields the final result
@@
Note that if the elasticity parameter is constant along income distribution ( $\alpha(p)=\alpha$), the generalized decomposition reduces to the usual Hall-Jones decomposition.